An amount of $200,000 is borrowed for 5 years at a rate of 12%. Make an amortization schedule showing the quarterly payment, the quarterly interest on the outstanding balance, the portion of the payment going toward reducing the debt, and the balance.
Answer #1
Step-1:Calculation of quarterly payment
Quarterly payment = Loan amount / Present value of annuity of 1
= $ 2,00,000.00 / 14.87747486
= $     13,443.14
Working:
Present value of annuity of 1 = (1-(1+i)^-n)/i Where,
= (1-(1+0.03)^-20)/0.03 i = 12%/4 = 0.03
= 14.87747486 n = 5*4 = 20
Step-2:Calculation of amortization schedule
Quarter Beginning Loan Interest Expense Quarterly Payment Reduction of debt Ending Loan
a b=a*0.03 c d=c-b e=a-d
1 $ 2,00,000.00 $       6,000.00 $     13,443.14 $       7,443.14 $ 1,92,556.86
2 $ 1,92,556.86 $       5,776.71 $     13,443.14 $       7,666.44 $ 1,84,890.42
3 $ 1,84,890.42 $       5,546.71 $     13,443.14 $       7,896.43 $ 1,76,993.99
4 $ 1,76,993.99 $       5,309.82 $     13,443.14 $       8,133.32 $ 1,68,860.67
5 $ 1,68,860.67 $       5,065.82 $     13,443.14 $       8,377.32 $ 1,60,483.35
6 $ 1,60,483.35 $       4,814.50 $     13,443.14 $       8,628.64 $ 1,51,854.71
7 $ 1,51,854.71 $       4,555.64 $     13,443.14 $       8,887.50 $ 1,42,967.21
8 $ 1,42,967.21 $       4,289.02 $     13,443.14 $       9,154.13 $ 1,33,813.08
9 $ 1,33,813.08 $       4,014.39 $     13,443.14 $       9,428.75 $ 1,24,384.34
10 $ 1,24,384.34 $       3,731.53 $     13,443.14 $       9,711.61 $ 1,14,672.72
11 $ 1,14,672.72 $       3,440.18 $     13,443.14 $     10,002.96 $ 1,04,669.76
12 $ 1,04,669.76 $       3,140.09 $     13,443.14 $     10,303.05 $     94,366.72
13 $     94,366.72 $       2,831.00 $     13,443.14 $     10,612.14 $     83,754.58
14 $     83,754.58 $       2,512.64 $     13,443.14 $     10,930.50 $     72,824.07
15 $     72,824.07 $       2,184.72 $     13,443.14 $     11,258.42 $     61,565.65
16 $     61,565.65 $       1,846.97 $     13,443.14 $     11,596.17 $     49,969.48
17 $     49,969.48 $       1,499.08 $     13,443.14 $     11,944.06 $     38,025.42
18 $     38,025.42 $       1,140.76 $     13,443.14 $     12,302.38 $     25,723.04
19 $     25,723.04 $           771.69 $     13,443.14 $     12,671.45 $     13,051.59
20 $     13,051.59 $           391.55 $     13,443.14 $     13,051.59 $              -0.00
Total $     68,862.83 $ 2,68,862.83 $ 3,37,725.66